# Multi-wavelength observations of the Luminous Fast Blue Optical Transient AT 2023fhn

### v1

@@ 4.1 UV-optical
We now compare the UV-optical constraints on AT 2023fhn’s light-curve with previous LFBOTs. All times used in Section 4 are in the rest-frames of the LFBOTs. Comparison data are corrected for Galactic extinction of =0.08 (AT 2018cow, Prentice et al., 2018) and =0.07 (ZTF 20acigmel, Perley et al., 2021), their UV light-curves (in absolute magnitude) are compared with AT 2023fhn in Figure 6. We fit the light-curve of AT 2018cow in 2 phases, early ( d) and late-time, with a fit of the form . For the fit to AT 2018cow, we assume that the late-time UV is dominated by residual transient emission (Sun et al., 2022, 2023; Chen et al., 2023; Inkenhaag et al., 2023). We shift the AT 2018cow best-fit up in absolute magnitude such that it lies between the early-time ATLAS -band and FORS2 -band AT 2023fhn detections (Ho et al., 2023b). The extrapolated curve passes below the late-time HST F225W and F336W upper limits reported in this work. Another LFBOT with good UV photometric coverage is ZTF 20acigmel, but here we consider only the early, pre-break phase due to a lack of late-time constraints. ZTF 20acigmel starts brighter than AT 2018cow and fades faster, whereas AT 2023fhn is the most luminous LFBOT yet at UV-optical wavelengths. A final addition to Figure 6 are bands of constant UV absolute magnitude, corresponding to late-time emission from black holes of different masses in the tidal disruption event model of Mummery et al. (2024). This model yielded a black hole mass of 103 M⊙ for AT 2018cow. Assuming similar evolution, the HST F336W point source upper limit for AT 2023fhn tentatively constrains the accreting black hole mass in a TDE interpretation to M⊙.
![Refer to caption](x3.png)
@@ 4.2 X-ray
Figure 7 shows our X-ray observations of AT 2023fhn, and the X-ray light-curves of other LFBOTs. The AT 2018cow broken power-law and late-time plateau fit of Migliori et al. (2024) is also shown. AT 2023fhn is the faintest LFBOT in X-rays at early times. Assuming a shallow decay initially, similar to AT 2022tsd, ZTF 20acigmel and AT 2018cow, the break time can be - at the latest - similar to AT 2018cow and ZTF 20acigmel. There appears to be a correlation between break time and X-ray luminosity, with brighter LFBOTs transitioning to a steeper decay at later times. Assuming instead that epochs 1 and 2 are on the same phase of the light-curve, the decay index (where ). Expectations for the X-ray decay rate are (shock power), (magnetar central engine) and (fallback, i.e. a TDE). Overall, the detections and upper-limits are consistent with AT 2023fhn behaving like a fainter version of previous LFBOTs in the X-ray band, and demonstrates that they can exhibit several orders of magnitude of variety in their X-ray luminosity.
![Refer to caption](extracted/5678772/xray_lc.png)
@@ 4.4 Radio
Using the scipy curve_fit function, and working with rest-frame times and central frequencies throughout this section, we fit equation 1 to the 90 day and 138 day (observer frame) data. At 90 days we have 6 data points (5 detections, 1 upper limit, we combine the 87 and 95 day data for this epoch), and at 138 days we have 7 data points (6 detections, 1 upper limit). There are 5 parameters to fit: , , , and . The best-fit values for these parameters and their uncertainties are listed in Table 7. The optically-thin spectral index of -0.59 (138 days) yields an electron energy spectral index of 2.2, which is relatively shallow - 2.5 is expected from theory, while values closer to 3 are often measured in gamma-ray bursts, tidal disruption events and supernovae (e.g. Chevalier & Fransson, 2006; Cendes et al., 2023). Values from other LFBOTs are also in the range 2–3 (Margutti et al., 2019; Ho et al., 2020; Coppejans et al., 2020; Yao et al., 2022; Bright et al., 2022).
@@
![Refer to caption](extracted/5678772/radio1.png)
![Refer to caption](extracted/5678772/radio2.png)
![Refer to caption](extracted/5678772/radio_fit.png)
![Refer to caption](extracted/5678772/blastwave.png)
![Refer to caption](x5.png)
@@ 5 Discussion
The host of AT 2023fhn, taking the spiral and satellite as one interacting system, is broadly consistent with the host galaxies of core-collapse supernovae, and slightly above four other LFBOT hosts in terms of specific star formation rate (although below the host of ZTF 18abvkwla). The host offset and faint, diffuse emission at the transient location (see Section 3.1) are consistent with the tail of the core-collapse supernova distribution (see also Chrimes et al., 2024). While the local and broader environment is consistent with a core-collapse origin, it is interesting to consider whether the high sSFR is related to tidal interaction between the spiral and satellite galaxy. Such interactions may be associated with an elevated tidal disrupted event (TDE) rate, which show a bias towards occurrence in post-starburst galaxies and galaxies undergoing interactions/mergers (French et al., 2016). The non-nuclear location of AT 2023fhn - at high offset from both the spiral and satellite - is difficult to explain in such a scenario (Chrimes et al., 2024). On the other hand, the measured optical magnitudes at the location of AT 2023fhn (see Table 4) allow a contribution from a point source with absolute magnitude as bright as -14. A significant contribution from a point source at the location of AT 2023fhn is disfavoured (as the precise location has similar brightness to its immediate surroundings, as explained in Section 3.1), but the presence of a globular cluster or ultra-compact dwarf galaxy - which may host massive black holes (e.g. Seth et al., 2014) - cannot be ruled out. The presence of such an undetected cluster or ultra compact dwarf galaxy would be consistent with the upper limit on the black hole mass inferred from our late-time UV observations of M⊙ (Mummery et al., 2024), given black hole - host galaxy/cluster mass relations (Kormendy & Ho, 2013; Lützgendorf et al., 2013).
@@ 6 Conclusions
An IMBH TDE interpretation remains possible, only if there exists a pre-existing dense CSM or if the early stages of the tidal disruption produce such an environment. The host galaxy is likely undergoing tidal interactions, which may elevate the TDE rate. While the non-nuclear location and host properties rate favour a core-collapse origin, the explosion of AT 2023fhn in an undetected globular cluster or ultra-compact dwarf galaxy cannot be ruled out.
@@ Acknowledgements.
A.A.C. acknowledges support from the European Space Agency (ESA) as an ESA Research Fellow. P.G.J. has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101095973). P.J.G. is partly supported by NRF SARChI Grant 111692.Observations analysed in this work were taken by the NASA/ESA Hubble Space Telescope under program 17238. This research has made use of software provided by the Chandra X-ray Center (CXC) in the application of the CIAO package (Fruscione et al., 2006). The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.

### v2

@@ 4.1 UV-optical
We now compare the UV-optical constraints on AT 2023fhn’s light-curve with previous LFBOTs. All times used in Section 4 are in the rest-frames of the LFBOTs. Comparison data are corrected for Galactic extinction of =0.08 (AT 2018cow, Prentice et al., 2018) and =0.07 (ZTF 20acigmel, Perley et al., 2021), their UV light-curves (in absolute magnitude) are compared with AT 2023fhn in Figure 6. We fit the light-curve of AT 2018cow in 2 phases, early ( d) and late-time, with a fit of the form . For the fit to AT 2018cow, we assume that the late-time UV is dominated by residual transient emission (Sun et al., 2022, 2023; Chen et al., 2023; Inkenhaag et al., 2023). We shift the AT 2018cow best-fit up in absolute magnitude such that it lies between the early-time ATLAS -band and FORS2 -band AT 2023fhn detections (Ho et al., 2023b). The extrapolated curve passes below the late-time HST F225W and F336W upper limits reported in this work. Another LFBOT with good UV photometric coverage is ZTF 20acigmel, but here we consider only the early, pre-break phase due to a lack of late-time constraints. ZTF 20acigmel starts brighter than AT 2018cow and fades faster, whereas AT 2023fhn is the most luminous LFBOT yet at UV-optical wavelengths. A final addition to Figure 6 are bands of constant UV absolute magnitude, corresponding to late-time emission from black holes of different masses in the tidal disruption event model of Mummery et al. (2024). This model yielded a black hole mass of 103 M⊙ for AT 2018cow. Assuming that AT 2023fhn had a similar light-curve shape to AT 2018cow - which may not be the case for accretion events around black holes of different masses - the HST F336W point source upper limit for AT 2023fhn tentatively constrains the accreting black hole mass in a TDE interpretation to M⊙.
@@ 4.2 X-ray
Figure 7 shows our X-ray observations of AT 2023fhn, and the X-ray light-curves of other LFBOTs. The AT 2018cow broken power-law and late-time plateau fit of Migliori et al. (2024) is also shown. AT 2023fhn is the faintest LFBOT in X-rays at early times. Assuming a shallow decay initially, similar to AT 2022tsd, ZTF 20acigmel and AT 2018cow, the break time can be - at the latest - similar to AT 2018cow and ZTF 20acigmel. Notwithstanding the small sample size, among the three LFBOTs with a clearly observed break in the X-ray light-curve (AT2018 cow, ZTF20 acigmel and AT 2022tsd), brighter LFBOTs seem to transition to a steeper decay at later times. Assuming instead that epochs 1 and 2 are on the same phase of the light-curve, the decay index (where ). Expectations for the X-ray decay rate are (shock power), (magnetar central engine) and (fallback, i.e. a TDE). Overall, the detections and upper-limits are consistent with AT 2023fhn behaving like a fainter version of previous LFBOTs in the X-ray band, and demonstrates that they can exhibit several orders of magnitude of variety in their X-ray luminosity.
@@ 4.4 Radio
Using the scipy curve_fit function, and working with rest-frame times and central frequencies throughout this section, we fit equation 1 to the 90 day and 138 day (observer frame) data. At 90 days we have 6 data points (5 detections, 1 upper limit, we combine the 87 and 95 day data for this epoch), and at 138 days we have 7 data points (6 detections, 1 upper limit). The data points (including limits, which use the rms as ) are given a 1/ weighting (with sigma_absolute = false). There are 5 parameters to fit: , , , and . The best-fit values for these parameters and their uncertainties (calculated from the covariance matrix output by curve_fit) are listed in Table 7. The optically-thin spectral index of -0.6 (138 days) yields an electron energy spectral index of 2.2, which is relatively shallow; 2.5 is expected from theory, while values closer to 3 are often measured in gamma-ray bursts, tidal disruption events and supernovae (e.g. Chevalier & Fransson, 2006; Cendes et al., 2024). Values from other LFBOTs are also in the range 2–3 (Margutti et al., 2019; Ho et al., 2020; Coppejans et al., 2020; Yao et al., 2022; Bright et al., 2022).
@@ 5 Discussion
The host of AT 2023fhn, taking the spiral and satellite as one interacting system, is broadly consistent with the host galaxies of core-collapse supernovae, and slightly above four other LFBOT hosts in terms of specific star formation rate (although below the host of ZTF 18abvkwla). The host offset and faint, diffuse emission at the transient location (see Section 3.1) are consistent with the tail of the core-collapse supernova distribution (see also Chrimes et al., 2024). While the local and broader environment is consistent with a core-collapse origin, it is interesting to consider whether the high sSFR is related to tidal interaction between the spiral and satellite galaxy. Such interactions may be associated with an elevated tidal disrupted event (TDE) rate, which show a bias towards occurrence in the nuclei of post-starburst galaxies and galaxies undergoing interactions/mergers (French et al., 2016). The non-nuclear location of AT 2023fhn - at high offset from both the spiral and satellite - is difficult to explain in such a scenario (Chrimes et al., 2024), since galaxy-galaxy interactions are unlikely to have much impact on the TDE rate in massive clusters well away from the nuclei of either galaxy. On the other hand, galaxy scale interactions may increase the chance of finding a massive (possibly IMBH-hosting, Fragione et al., 2018) cluster in such a location (e.g. Napolitano et al., 2022), even if they do not increase the TDE rate within the cluster.
The measured optical magnitudes at the location of AT 2023fhn (see Table 4) allow a contribution from a point source with absolute magnitude as bright as -14. A significant contribution from a point source at the location of AT 2023fhn is disfavoured (as the precise location has similar brightness to its immediate surroundings, as explained in Section 3.1), but the presence of a globular cluster or ultra-compact dwarf galaxy - which may host massive black holes (e.g. Seth et al., 2014) - cannot be ruled out. The presence of such an undetected cluster or ultra compact dwarf galaxy would be consistent with the upper limit on the black hole mass inferred from our late-time UV observations of M⊙ (Mummery et al., 2024), given black hole - host galaxy/cluster mass relations (Kormendy & Ho, 2013; Lützgendorf et al., 2013).
@@ 6 Conclusions
An IMBH TDE interpretation remains possible, only if there exists a pre-existing dense CSM or if the early stages of the tidal disruption produce such an environment. The host galaxy is likely undergoing tidal interactions, which may increase the chance of finding clusters (and hence TDEs) at high project offsets. While the non-nuclear location and host properties rate likely favour a core-collapse origin, the explosion of AT 2023fhn in an undetected globular cluster or ultra-compact dwarf galaxy cannot be ruled out.
@@ Acknowledgements.
We thank the anonymous referee for their constructive feedback on this manuscript.A.A.C. acknowledges support from the European Space Agency (ESA) as an ESA Research Fellow. P.G.J. has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101095973). P.J.G. is partly supported by NRF SARChI Grant 111692.
Observations analysed in this work were taken by the NASA/ESA Hubble Space Telescope under program 17238. This research has made use of software provided by the Chandra X-ray Center (CXC) in the application of the CIAO package (Fruscione et al., 2006). The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.